A taxi charges an initial charge of $3. Your fare will be computed as the $3 charge plus 20 cents for every mile traveled.
a) Write the cost (C) of a taxi ride in terms of the number of miles ridden (X).
b) If the fare is $4.70, how far did you ride?

Answer by ntnk(54)

To solve this problem, begin by multiplying both sides of the equation by the variable t to cancel it from the right move it to the left side:
20*t=60*t/t
20*t=60
Now, divide 20 from both sides of the equation:
20/20*t=60/20
t=3
Remember: When you have a number divided by itself, it becomes 1.

Check: 20=60/(3)

Solved By: NtNk